# KEN KANGDA WREN1

### v4

@@ 1.1 Usual-Case Assets and Information Structure
  7. 7.
[…]
1.2 The FTAP Baseline under Model-Risk
Translated to our setup, FTAP states: market $\big{(}Y,\{S_{t}\};\text{ }\mathcal{S}^{T}_{\mathcal{B}},\{\mathbf{F}_{t}\},\{\mathbf{F}_{T}\};\text{ }\pi^{B}_{0}\mathbf{W}_{B}^{T}\big{)}$, with $\{\mathbf{F}_{t}\}$-adapted asset-prices $\{S_{t}\},\text{ }S_{T}\equiv Y$, is viable iff. there is an equivalent measure $\widehat{\pi^{B}_{0}\mathbf{W}^{T}_{B}}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, the RNE measure, such that $S_{t}=\hat{\mathbf{E}}[Y\big{|}\mathbf{F}_{t}]$, $\forall t<T$, where RNE expectation $\hat{\mathbf{E}}[\cdot{|}\cdot]$ may be written as:
@@ 1.2 The FTAP Baseline under Model-Risk
with $\Sigma_{B}{(\cdot)^{B}}{(\cdot)^{B}}$ denoting convex combinations, such as the one under RNE beliefs $\hat{\pi}^{B}_{t}$, as calculated under Bayes’ Rule (3) given $\hat{\pi}^{B}_{0}$ and RNE log-LR $\hat{l}^{+/-}_{t}:=\log\frac{d\hat{\mathbf{W}}^{T}_{+}|_{t}}{d\hat{\mathbf{W}}^{T}_{-}|_{t}}$.
This market, defined by (1-6), is complete (Martingale Representation Theorem), and so the RNE measure, unique.
Viability by no means ensures economic validity (e.g. any pricing based on RNE measures of the form $\tilde{\pi}^{B}_{0}\tilde{\mathbf{W}}^{T}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ is blind to $B$-information); indeed, any pricing that is viable to any member of the equivalence class $[{\pi}^{B}_{0}\mathbf{W}_{B}^{T}]$ is viable to all in the class. That is, under FTAP-viability, ’anything goes’ (almost). What then characterises the processes (and the RNE measures) that do correspond to plausible economic theories? It turns out that having economically interpretable RNE price-of-model-risk is the key.
Consider risk-premium, the gap between ex-ante price and expectation: at any $t<T$, given reference model-risk belief $\{{\pi}^{B}_{t}\}$ ((3)),
$RP_{t}:=Y_{t}-S_{t}:=\Sigma_{B}{\pi}^{B}_{t}Y^{B}_{t}-S_{t},$ (7)
and its RNE decomposition (by inserting ±$\pm\Sigma_{B}\hat{\pi}^{B}_{t}{{Y}^{B}_{t}}$):
@@ 1.3 The Economic Decomposition of Risk-Pricing under Model-Risk
Regardless of the economic pricing scheme under model-risk, classical or otherwise, the resulting asset-price $\{S_{t}\}$, with risk-premia $\{RP_{t}\}$ and RNE formulation (6-11), admits an economic decomposition of the form (16-19), provided the existence of 1) some reference model-risk beliefs $\{{\pi}^{B}_{t}\}$ and 2) $B$-sure pricing $\{S^{B}_{t}\}$ under the same scheme when $B$-sure:
@@
$RP^{B}_{t}:=y_{B}(T-t,t)-\check{y}_{B}(T-t,t)=:RP^{B}(t)\geq 0,$ (13)
@@
$RP^{B}(t)=\lim_{\sigma^{\pi}_{t}\to 0}\hat{RP}^{B}(t),\text{ }\text{i.e. }S^{B}_{t}=\lim_{\sigma^{\pi}_{t}\to 0}\hat{S}^{B}_{t},$ (15)
where (14) ensures consistency with asset-property Item-7, Section 1.1, and (15), internal consistency of the pricing scheme under model-risk.
Remark 5.
Usual-case model-sure pricing is simple: static noise discounted by a deterministic drift. With model-risk $B$, the only new input to economic pricing (recursive optimisation taking drivers of the expected preference-value as inputs) is the reference model-risk belief $\{{\pi}^{B}_{t}\}$.
At any $t<T$, viable pricing must satisfy $S_{t}\in(S^{-}_{t},S^{+}_{t})$, so there always exists a well-defined pricing coefficient ${A}^{B}_{t}\in(0,1)$ such that:
@@
where the economic price-of-$B$-risk $\{k^{A}_{t}\}$ should be free from explicit dependence on $B$-sure pricing, the latter being based on decisions in the absence of $B$-risk.
As such, once obtained, price-of-$B$-risk $\{k^{A}_{t}\}$ may generate the coefficients $\{A^{B}_{t}\}$ via (19) for asset-pricing under $B$-risk, given consistent but otherwise arbitrary $B$-sure prices $\{{S^{\prime}}^{B}_{t}\}:=\{Y^{B}_{t}\}-\{RP^{\prime B}(t)\}$, thus separating the pricing of model and non-model risks.
This is attractive if for a useful range of risk-premia, the prices so constructed:
$\{S^{\prime}_{t}\}\equiv\{S^{(RP^{\prime}|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}{S^{\prime}}^{B}_{t}\}=\{S^{(0|A)}_{t}\}-\text{ }\{\Sigma_{B}A^{B}_{t}{RP^{\prime}}^{B}(t)\},\text{ with }$ (20)
$\{S^{(0|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}Y^{B}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}(t)\}=\{Y_{t}\}-\text{ }\{k^{A}_{t}\sigma^{\pi}_{t}\cdot{Y}^{\Delta}(t)\},$ (21)
are viable; both are one way or another the sum of a martingale and a drift. Observe that all applied economic asset-pricing theories, continuous and regular near risk-neutrality, are already viable written in above form, at least for sufficiently small risk-discounts.
@@
In ambiguity-aversion settings, the primary utility function, enforcing risk-aversion, can be dialled towards near neutrality (linearity) while secondary utility/variational-control continues to exert uncertainty discounts. In classical settings, including parameter-learning, where a single utility must be optimised, the same may be achieved by replacing the payoff variable with its up-to-date model-sure expectation perturbed by a small enough noise.
The real question is then the range of viability for processes (20-21), and most relevant to us, if they have convenient RNE formulations. For this, we must characterise the economic price-of-$B$-risk $\{k^{A}_{t}\}$ ((19)) more fully. Its origin suggests the following.
@@
Asset-pricing under model-risk $B$ is said to be normal if its price-of-model-risk $\{k^{A}_{t}\}$ at any $t\ll T$: 1) depends deterministically on reference model-risk belief $\{\pi^{B}_{t}\}$ in a time-homogenous manner [^2] Economic risk-pricing, solution to some optimisation-based pricing scheme, is measurable to the natural filtration of its input, ($(\{Z_{t}\},\{\pi^{B}_{t}(\mathbf{D}_{t})\})$, and so a function of its level to date (Proposition 4.9 of Breiman (1992)). Any explicit dependence on time $t$ or $Z_{t}$ itself, if at all, is presumed weak or non-existent, at least for $t\ll T$. such that the dependence is twice-differentiable and non-negative; 2) satisfies $\lim_{\sigma^{\pi}_{t}\to 0}k^{A}_{t}=0$; 3) generates viable pricing via (20-21) under an economically relevant range of risk-premium choices.
@@
Property-1&2 are shared by most if not all existing model-risk pricing theories. Property-3 will be shown to hold for usual-case assets under model-risk.
@@ 2 Canonical RNE Asset-Pricing under Model-Risk
2.1 Viable Normal Model-Risk Pricing and its RNE measure
@@
For usual-case assets under model-risk $B$, with reference model-risk process $\{\pi^{B}_{t}\}$ based on some reference belief $\pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$ from the set of model-uncertainty, normal pricing (Definition 1) is viable iff. it has this canonical form: at any $t<T$,
@@
where model-risk only pricing $\{S^{(0|{\Pi})}_{t}\}=\{\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}\}$ ((21)) is given by some RNE model-risk inference process $\{\Pi^{B}_{t}\}$ based on a RNE measure of the form $\Pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
Proof. See Appendix B. Briefly, any viable model-risk only pricing $\{S^{(0|A)}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}(t)\}$ must have drift $\{Y^{\Delta}(t)dA^{+}_{t}\}$ offset $\{dY^{-}_{t}+A^{+}_{t}dY^{\Delta}(t)\}$ in expectation under some RNE measure, possible iff. it belongs to $\mathcal{L}_{\mathcal{S}}$, under which the two drift separately vanish in expectation, due to $\{A^{+}_{t}\}$ being a twice-differentiable function of model-risk inference $\{\pi^{+}_{t}\}$ driven partly/wholly by $\{dY^{-}_{t}\}$. The viability of (22), as a martingale plus a well-behaved drift, follows.
Canonical pricing thus allows the pricing of model and non-model risks to have separate RNE formulations that are each easy to identify and apply. Note however that the overall RNE measure of viable normal asset-pricing (22) under model-risk $B$, although guaranteed to exist, remains unidentifiable in general, even for simple cases.
@@
Note that canonical model-risk only pricing $\{S^{(0|{\Pi})}_{t}\}$ has identical RNE and economic price-of-model-risk ((11) and (19)). This indeed exhausts the set of price processes with this property (Lemma 1 and Remark 14, Appendix A). In fact any process with non-trivial risk-pricing of the non-model risks (i.e. $\{RP^{B}_{t}\}\not=0$) must have an RNE price-of-model-risk that is not a time-homogeneous deterministic function of any reference model-risk belief based on the set $\mathcal{L}_{\mathcal{S}}$ of model-uncertainty [^3] Any viable price process whose RNE measure is in $\mathcal{L}_{\mathcal{S}}$ must have the form $\{\Sigma_{B}A^{B}_{t}S^{B}_{t}\}=\{\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}\}$, so $\forall t<T$, $\lim_{\sigma^{\pi}_{t}\to 0}S_{t}=Y^{B}_{t}$ since $\lim_{\sigma^{\pi}_{t}\to 0}\sigma^{\Pi}_{t}=0$, but $\lim_{\sigma^{\pi}_{t}\to 0}{S}_{t}=S^{B}_{t}=Y^{B}_{t}-RP^{B}_{t}$ given (15); thus $\{RP^{B}_{t}\}=0$.. This echos the observation that parameter-learning makes RNE formulations path-dependent (Guidolin and Timmermann (2007)).
2.2 Properties of Canonical Model-Risk Pricing
2.2.1 The 1-Parameter Price-of-Model-Risk
Corollary 1.
Canonical model-risk premium $\{B\_RP_{t}\}$ and price-of-model-risk $\{k^{\Pi}_{t}\}$ are as follows:
$B\_RP_{t}$ $=({\pi}^{+}_{t}-{\Pi}^{+}_{t}){S}^{\Delta}(t),\text{ }t<T,$ (23)
$k^{\Pi}_{t}$ $=(K^{\frac{1}{2}}-K^{-\frac{1}{2}})\sigma^{\Pi}_{t},\text{ }t<T,$ (24)
with $\sigma^{\Pi}_{t}:=({\Pi}^{+}_{t}\underline{\Pi}^{+}_{t})^{\frac{1}{2}}$, and $K:=\frac{O_{f}[{\pi}^{+}_{0}]}{O_{f}[{\Pi}^{+}_{0}]}=\frac{O_{f}[{\pi}^{+}_{t}]}{O_{f}[{\Pi}^{+}_{t}]}$, a conserved quantity of risk-inference ((3)).
@@ 2.2.1 The 1-Parameter Price-of-Model-Risk
[…]
Remark 10.
The sign of the difference process $\{{\pi}^{+}_{t}\}-\{{\Pi}^{+}_{t}\}$ is fixed ((3)); so unless model-risk impact is consistent (Item-7, Section 1.1), model-risk premia $\{B\_RP_{t}\}$ must turn negative at times: inconsistent model-risks cannot have interpretable viable economic risk-premia.
2.2.2 Connection with Familiar Price Dynamics
[…]
$S_{t}=S^{(\check{rp}|\Pi)}_{t}=S^{(0|\Pi)}_{t}-\check{rp}(t)=Y^{-}_{t}-\check{rp}(t)+S^{\Delta}(t)\Pi^{+}_{t},$ (25)
$dS_{t}(B)=dY^{-}_{t}(B)+\check{R}_{t}dt-\Pi^{+}_{j}(B)r^{Z\Delta}_{t}dt+S^{\Delta}(t)d\Pi^{+}_{t}(B).$ (26)
Given Section 1.1 and inference dynamics (A.2-A.3), we have $\{d\Pi^{+}_{t}\}$ as below, with two sources of noise, one from $\{Z_{t}\}$, given by $\{\sigma^{lZ}_{t}dw^{Z}_{t}\}$, with $\sigma^{lZ}_{t}:=r^{Z\Delta}_{t}/\sigma^{Z}_{t}$ (’signal-to-noise’ ratio), one from independent $B$-informative data $\{D_{t}\}$, modelled in parallel ((2)), which can be absorbed via $\{\sigma^{l}_{t}dw^{\prime}_{t}\}=\{\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}\}$, with $(\sigma^{l}_{t})^{2}=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$:
$d\Pi^{+}_{t}(B)=\mu^{\Pi,B}_{t}+\sigma^{\Pi,l}_{t}dw^{\prime}_{t},\text{ with},$ (27)
$\mu^{\Pi,B}_{t}:=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\Big{(}{{\mathbf{1}_{+}}}(B)-{\Pi}^{+}_{t}(B)\Big{)},$ (28)

### v6

@@ 1.2 The FTAP Baseline under Model-Risk
Such FTAP-viable asset-pricing under model-risk $B$ may be written always as follows:
@@
where $\Sigma_{B}{(\cdot)^{B}}{(\cdot)^{B}}$ denotes averages weighted by $B$-outcomes, such as by RNE beliefs $\hat{\pi}^{B}_{t}$ under Bayes’ Rule (3) given $\hat{\pi}^{B}_{0}$ and RNE log-LR $\hat{l}^{+/-}_{t}:=\log\frac{d\hat{\mathbf{W}}^{T}_{+}|_{t}}{d\hat{\mathbf{W}}^{T}_{-}|_{t}}$.
Viability by no means ensures economic validity (e.g. any pricing based on RNE measures of the form $\tilde{\pi}^{B}_{0}\tilde{\mathbf{W}}^{T}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ is blind to $B$-informative data); indeed, any pricing that is viable to any member of the equivalence class ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}]$ is viable to all in the class. In this sense, ’anything goes’ under FTAP-viability. What characterises the processes and the RNE measures that do correspond to plausible economic theories?
For answers, consider risk-premium, the gap between ex-ante pricing and expectation, non-negative under economic asset-pricing: at $\forall t<T$, given reference model-risk belief $\{{\pi}^{B}_{t}\}$,
$RP_{t}:=Y_{t}-S_{t}=\Sigma_{B}{\pi}^{B}_{t}Y^{B}_{t}-S_{t}\geq 0.$ (7)
It has RNE decomposition (by inserting $\pm\Sigma_{B}\hat{\pi}^{B}_{t}{{Y}^{B}_{t}}$):
@@
Decomposition (8-11) exists whenever the RNE measure exists, that is, for all viable pricing schemes. However, its elements, $\{\hat{S}^{B}_{t}\}$ and $\{{k}^{\hat{\pi}}_{t}\}$, need not correspond to economically interpretable entities. It turns out that establishing such a correspondence, albeit indirectly, is the key to finding economically valid asset-prices (out of mere viable ones).
@@ 1.3 The Economic Decomposition of Risk-Pricing under Model-Risk
Just as any viable price process admits a RNE decomposition of the embedded risk-premia, any economic pricing scheme under model-risk, classical or otherwise, with asset-price $\{S_{t}\}$, risk-premia $\{RP_{t}\}$ and RNE formulation (6-11), admits an economic decomposition, (16-19). It is unique and well-defined provided the existence of 1) a reference model-risk belief process $\{{\pi}^{B}_{t}\}$; 2) viable $B$-sure pricing $\{S^{B}_{t}\}$ produced under the same scheme when $B$-sure:
@@
$RP^{B}_{t}:=Y^{B}_{t}-S^{B}_{t}=y_{B}(T-t,t)-\check{y}_{B}(T-t,t)=:RP^{B}(t)\geq 0,$ (13)
where $\check{\mathbf{E}}_{{B}}[\cdot|\cdot]$ denotes expectations under its $B$-sure RNE measure $\check{\mathbf{W}}^{T}_{B}\sim{\mathbf{W}}^{T}_{B}$, leading to RNE outlook $\check{y}_{B}(T-t,t)=\int_{t}^{T}\check{r}_{B,s}ds$ as in (4), and $B$-sure risk-premium $RP^{B}(t)$ satisfying: $\forall t\leq T$,
$S^{B}_{t}=\lim_{\sigma^{\pi}_{t}\to 0}\hat{S}^{B}_{t},\text{ }\text{i.e. }RP^{B}(t)=\lim_{\sigma^{\pi}_{t}\to 0}\hat{RP}^{B}(t),$ (15)
where (14) ensures consistency with asset-property Item-2, Section 1.1, while (15), the internal consistency of the pricing scheme under model-risk $B$.
Remark 4.
Model-sure usual-case pricing is simple: static noise compensated by a predictable drift. Adding model-risk, the only new input to economic pricing (recursive optimisation taking drivers of the expected preference-value as inputs) is the reference model-risk belief $\{{\pi}^{B}_{t}\}$.
@@
which parallels RNE decomposition (8-11), but pricing coefficients $\{A^{B}_{t}\}\in(0,1)$, unlike $\{\hat{\pi}^{B}_{t}\}$, need not be a conditional-probability process. We show shortly that they in fact must be.
The economic price-of-$B$-risk $\{k^{A}_{t}\}$ should have no explicit dependencies on $B$-sure prices, the latter being decisions in the absence of $B$-risk. Thus, once devised, process $\{k^{A}_{t}\}$ can generate asset-prices under $B$-risk via (19) and (16) given any consistent but otherwise arbitrary $B$-sure pricing $\{{S}^{B}_{t}\}\equiv\{Y^{B}_{t}\}-\{RP^{B}(t)\}$ ((13)), thus separating the pricing of model and non-model risks. This is attractive if for a useful range of risk-discount choices, the asset-pricing so achieved,
$\{S^{(RP|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}{S}^{B}_{t}\}=\{S^{(0|A)}_{t}\}\text{ }-\text{ }\{\Sigma_{B}A^{B}_{t}{RP}^{B}(t)\},\text{ with }$ (20)
$\{S^{(0|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}Y^{B}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}(t)\}=\{Y_{t}\}\text{ }-\text{ }\{k^{A}_{t}\sigma^{\pi}_{t}\cdot{Y}^{\Delta}(t)\},$ (21)
is viable. Observe that both above can be the sum of a martingale and a well-behaved drift, and that all applied economic pricing theories, continuous and regular near risk-neutrality, are already viable when written in this form at least for small enough risk-discounts.
@@
In ambiguity-aversion settings, the primary utility function, enforcing risk-aversion, and the secondary utility/variational-control, enforcing uncertainty-aversion, can be dialled independently and arbitrarily close to neutrality (linearity), to achieve the above separation. In classical and some parameter-learning settings, with a single utility, the above obtains by replacing payoffs with their model-sure expectations perturbed by suitably small noise.
The real question is then the range of viability for (20-21) and if they have tractable RNE formulations. For this, let us characterise the economic price-of-$B$-risk $\{k^{A}_{t}\}$ better:
@@
Asset-pricing under model-risk $B$ is said to be usual if its price-of-$B$-risk $k^{A}_{t}$, (19), is a time-homogenous twice-differentiable function of its model-risk belief [^3] Economic risk-pricing, as the solution to some optimisation scheme, is measurable to the natural filtration of its input $(\{Z_{t}\},\{\pi^{B}_{t}(\mathbf{D}_{t})\})$; so it is at any time $t$ a function of $(Z_{t},\pi^{B}_{t}(\mathbf{D}_{t}))$ (Proposition 4.9 of Breiman (1992)). Any explicit dependence on $t$ or $Z_{t}$, if at all, is presumed here weak or non-existent, at least for $t\ll T$. $\pi^{B}_{t}$, $\forall t\ll T$.
@@
This makes $\{k^{A}_{t}\}$ a time-homogeneous Ito process of risk-inference $\{\pi^{B}_{t}\}$, a common feature of economic risk-pricing theories: priced risks are captured as state-variables so that the resulting risk-pricing is their function and hence path-independent with respect to them. It will be seen (via Proposition 1 and Corollary 1) that this is responsible for all the ’nice’ properties of $\{k^{A}_{t}\}$: fixed-signed, with $\lim_{\sigma^{\pi}_{t}\to 0}k^{A}_{t}=0$ (ensuring (15)), and generating prices (20-21) that follow familiar dynamics and allow a wide range of viable risk-premium choices.
@@ 2 Canonical RNE Asset-Pricing under Model-Risk
2.1 Viable, Usual, Model-Risk Pricing
@@
Under model-risk $B$ and model-risk inference $\{\pi^{B}_{t}\}$ based on a reference belief $\pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$ that belongs to the set $\mathcal{L}_{\mathcal{S}}$ of model-uncertainty (Remark 2), any usual pricing (Definition 1) of any usual-case asset is viable iff. it has the canonical form below: $\forall t<T$,
@@
where model-risk only asset-pricing $S^{(0|{\Pi})}_{t}=\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}$ (see (21)) is given by a RNE model-risk inference process $\{\Pi^{B}_{t}\}$ under a RNE measure of the form $\Pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
Proof. Appendix B. Briefly, viable pricing $S^{(0|A)}_{t}=Y^{-}_{t}+A^{+}_{t}Y^{\Delta}(t)$ must see $dY^{-}_{t}+A^{+}_{t}dY^{\Delta}(t)$ offset $Y^{\Delta}(t)dA^{+}_{t}$ in expectation under a RNE measure, possible iff. the two expectations vanish separately due to $\{A^{+}_{t}\}$ being an Ito process of $\{\pi^{+}_{t}\}$, thus limiting the RNE to the form stated. The viability of (22) then follows from the ’good behaviour’ of $\{\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t}\}$
Canonical model-risk only pricing $\{S^{(0|{\Pi})}_{t}\}$, with $\{RP^{B}_{t}\}=0$, has identical economic and RNE price-of-model-risk ((19) and (11)); indeed any pricing with this property must be model-risk only (with $\{RP^{B}_{t}\}=0$) [^4] Any such pricing must have its RNE measure in $\mathcal{L}_{\mathcal{S}}$ and so have the form $\{\Sigma_{B}A^{B}_{t}S^{B}_{t}\}=\{\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}\}$. Hence, $\forall t<T$, $\lim_{\sigma^{\pi}_{t}\to 0}S_{t}=Y^{B}_{t}$ since $\lim_{\sigma^{\pi}_{t}\to 0}\sigma^{\Pi}_{t}=0$, but $\lim_{\sigma^{\pi}_{t}\to 0}{S}_{t}=S^{B}_{t}=Y^{B}_{t}-RP^{B}_{t}$ given (15); thus $\{RP^{B}_{t}\}=0$.. Given Definition 1 and Remark 6, this means that viable prices (16) that have non-trivial $B$-sure risk-pricing (with $\{RP^{B}_{t}\}\not=0$) must have a RNE price-of-model-risk that is path-dependent with respect to any model-risk process $\{\pi^{B}_{t}\}$ derived from the set $\mathcal{L}_{\mathcal{S}}$ of model-uncertainty. This formalises the observation of Guidolin and Timmermann (2007) that learning makes RNE probabilities path-dependent and so intractable.
The source of this difficulty is basically (15) (see footnote-4), given Lemma 1 and the fact that the operation of $B$-conditioning and of risk-neutralising (reference beliefs such as $\pi^{B}_{0}\mathbf{W}^{T}_{B}$) do not commute in general. Canonical pricing (22) does not remove this intrinsic issue but avoids it by separating the RNE pricing of model and non-model risks, each by itself easy and tractable. The total RNE measure of (22), though guaranteed, remains intractable.
2.2 Properties of Canonical Model-Risk Pricing
2.2.1 The 1-Parameter Price-of-Model-Risk Formula
Corollary 1.
Canonical price-of-model-risk $\{k^{\Pi}_{t}\}:=\{\frac{{\pi}^{+}_{t}-\Pi^{+}_{t}}{\sigma^{\pi}_{t}}\}$ (see (19) or (11)) satisfies:
$k^{\Pi}_{t}=(K^{\frac{1}{2}}-K^{-\frac{1}{2}})\sigma^{\Pi}_{t},\text{ }\forall t<\infty,$ (23)
with $\sigma^{\Pi}_{t}:=({\Pi}^{+}_{t}\underline{\Pi}^{+}_{t})^{\frac{1}{2}}$ and $K:=\frac{O_{f}[{\pi}^{+}_{0}]}{O_{f}[{\Pi}^{+}_{0}]}=\frac{O_{f}[{\pi}^{+}_{t}]}{O_{f}[{\Pi}^{+}_{t}]}$, a conserved quantity.
@@ 2.2.1 The 1-Parameter Price-of-Model-Risk Formula
This follows from Bayes’ Rule (3). Price-of-model-risk $k^{\Pi}_{t}$ is positive iff. $K>1$. It takes value $k_{\frac{1}{2}}:=(K-1)/(K+1)$ at peak risk $\sigma^{\pi}_{t}=\frac{1}{2}$, where model-risk pricing offers a risk-premium of $\frac{1}{2}k_{\frac{1}{2}}{S}^{\Delta}(t)$ (see (18)) and a gain-to-loss ratio [^5] Gain-to-loss with respect to $B$-outcomes is always $O_{f}[{\Pi}^{+}_{t}]^{-1}$, but $O_{f}[{\pi}^{b}_{t}]=1$ at peak $B$-risk, hence the result. of $(1+k_{\frac{1}{2}})/(1-k_{\frac{1}{2}})\equiv K$. To any buyers (sellers), the higher (lower) $K$ is, the better; it is natural in theory and practice to set $K\in(1,2)$ for ’competitively priced’ model-risk, with $K-1\not\gg 1$ in any case.
@@
The difference process $\{{\pi}^{+}_{t}\}-\{{\Pi}^{+}_{t}\}=\{k^{\Pi}_{t}\sigma^{\pi}_{t}\}$ has a fixed sign. Unless the model-risk is consistent (see Item-2, Section 1.1, and (14)), viable model-risk premia $\{B\_RP_{t}\}$ (see (18)) must be negative sometimes and so correspond to no known economic risk-pricing theories.
2.2.2 Connection with Familiar Price-Dynamics
For better focus, let any $B$-sure risk-premia be $B$-independent henceforth, to be written as $\check{rp}(t)=\int_{t}^{T}\check{R}_{s}ds>0$, with $\check{R}_{t}$ denoting the $B$-sure risk-premium drift at any $t<T$; note then $Y^{\Delta}(t)=S^{\Delta}(t)=\int_{t}^{T}r^{Z\Delta}_{s}ds$ as a result. Canonical pricing (22) becomes:
$S_{t}=S^{(0|\Pi)}_{t}-\check{rp}(t)=Y^{-}_{t}+S^{\Delta}(t)\Pi^{+}_{t}-\check{rp}(t),$ (24)
$dS_{t}(B)=dY^{-}_{t}(B)+S^{\Delta}(t)d\Pi^{+}_{t}(B)-r^{Z\Delta}_{t}\Pi^{+}_{j}(B)dt+\check{R}_{t}dt.$ (25)
By setup (Section 1.1) and inference dynamics (A.2-A.3), process $\{d\Pi^{+}_{t}\}$ has two sources of noise, one from $\{Z_{t}\}$, given by $\{\sigma^{lZ}_{t}dw^{Z}_{t}\}$, with $\sigma^{lZ}_{t}:=r^{Z\Delta}_{t}/\sigma^{Z}_{t}\ll 1$ (’signal-to-noise’), one from independent $B$-informative data $\{D_{t}\}$, modelled in parallel ((2)), which can be absorbed through $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$ and $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$; then:
$d\Pi^{+}_{t}(B)=\mu^{\Pi,B}_{t}dt+\sigma^{\Pi,l}_{t}dw_{t},\text{ with},$ (26)
$\mu^{\Pi,B}_{t}:=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot\big{(}{{\mathbf{1}_{+}}}(B)-{\Pi}^{+}_{t}(B)\big{)},$ (27)